# Liouville Flow Importance Sampler

### v1

@@ D.10 Sequential Monte Carlo
Sequential Monte Carlo (Del Moral et al., 2006) serves both as the reference method as well as the gold standard for those problems that are not analytically tractable. A series of similar studies including AFT (Arbel et al., 2021a), PIS (Zhang & Chen, 2022), and DDS (Vargas et al., 2023a) took the same approach. We found that the SMC implementations in AFT and DDS were unnecessarily complicated. In particular, the leapfrog step size in the Hamiltonian Monte Carlo (HMC) kernel in those implementations depends on the progression of the tempered scales. Although progression-dependent step size was meticulously tuned, we found that a fixed step size with a sufficiently long HMC kernel and sufficiently high Effective Sample Size (ESS) threshold for triggering resampling can achieve similar results. In addition, the code base in Arbel et al. (2021b) also contains an error that the number of HMC kernels per Monte Carlo step (hmc_steps_per_iter specified in the configuration files) was not properly performed.
@@ Appendix E Ablation study of Annealed Flow Transport Monte Carlo Sampler
We tested the AFTMC sampler (with 256 scales) without the MC kernel on the Gaussian Mixture, Funnel, Variational Auto-Encoder, and Log-Gaussian Cox Process problems. We juxtapose the results of (1) SMC with $T=1024$ as the gold standard, (2) SMC with $T=256$ as a reference, (3) AFT with a Monte Carlo kernel and $T=256$, and (4) AFT without the Monte Carlo kernel and $T=256$ in Table 14. The results indicate that the MC kernel, instead of the flow-transport kernel, is the major functioning component in estimating the marginal likelihood $\log\mathcal{Z}$. The AFTMC algorithm with only the MC kernel is functionally identical to SMC. The numerical experiments on these datasets seemed to suggest that the flow transport implemented as a normalizing flow, parametrized by variational inference, does not seem to function as expected.

### v2

@@ D.10 Sequential Monte Carlo
Sequential Monte Carlo (Del Moral et al., 2006) serves both as the reference method as well as the gold standard for those problems that are not analytically tractable. A series of similar studies including AFT (Arbel et al., 2021a), PIS (Zhang & Chen, 2022), and DDS (Vargas et al., 2023a) took the same approach. We found that the SMC implementations in AFT and DDS were unnecessarily complicated. In particular, the leapfrog step size in the Hamiltonian Monte Carlo (HMC) kernel in those implementations depends on the progression of the tempered scales. Although progression-dependent step size was meticulously tuned, we found that a fixed step size with a sufficiently long HMC kernel and sufficiently high Effective Sample Size (ESS) threshold for triggering resampling can achieve similar results.
@@ Appendix E Ablation study of Annealed Flow Transport Monte Carlo Sampler
We tested the AFTMC sampler (with 256 scales) without the MC kernel on the Gaussian Mixture, Funnel, Variational Auto-Encoder, and Log-Gaussian Cox Process problems. We juxtapose the results of (1) SMC with $T=1024$ as the gold standard, (2) SMC with $T=256$ as a reference, (3) AFT with a Monte Carlo kernel and $T=256$, and (4) AFT without the Monte Carlo kernel and $T=256$ in Table 14. The results indicate that the MC kernel, instead of the flow-transport kernel, is the major functioning component in estimating the marginal likelihood $\log\mathcal{Z}$. The AFTMC algorithm with only the MC kernel is functionally identical to SMC. The numerical experiments on these datasets suggest that the flow transport, implemented as a normalizing flow and parametrized by variational inference, does not function independently from the MCMC steps, which still need to be designed and fine-tuned akin to typical SMCs. This contrasts with the proposed LFIS, which solely depends on the flow for transporting samples without designing additional MCMC kernels.